OpenStudy (anonymous):

I need help with 9x^2+9x-4=0. I need solutions. Someone please help me???????

OpenStudy (anonymous):

is it -2?

OpenStudy (anonymous):

9x^2+9x-4=0 9x^2+12x-3x-4=0 3x(3x-4-1(3x-4)=0 (3x-4)(3x-1)=0 x=4/3,1/3.

OpenStudy (anonymous):

so -4/3 and 1/3

OpenStudy (anonymous):

oops...yes -4/3 and 1/3 :)